(Average

expected return and

risk​)

Given the​ holding-period returns shown​ here, calculate the average returns and the standard deviations for the Kaifu Corporation and for the market.

MONTH

KAIFU CORP.

MARKET

1

4​%

2​%

2

7​%

  

2​%

3

0​%

  

0​%

4

1​%

−1​%

​(Click

on the icon located on the​ top-right corner of the data table above in order to copy its contents into a

spreadsheet.​)

a. The average monthly return for Kaifu Corporation is

​%.

​(Round to two decimal​ places.) The average monthly return for the market is

​%.

​(Round to two decimal​ places.) b. The standard deviation for Kaifu Corporation is

​%.

​(Round to two decimal​ places.) The standard deviation for the market is

​%.

​(Round to two decimal​ places.)
Answer #1

Average return for stock=(4%+7%+0%+1%)/4=3%

Standard Deviation for stock=SQRT((4%-3%)^2+(7%-3%)^2+(0%-3%)^2+(1%-3%)^2)/SQRT(3)=3.1622777%

Average return for market=(2%+2%+0%-1%)/4=0.75%

Standard Deviation for market=SQRT((2%-0.75%)^2+(2%-0.75%)^2+(0%-0.75%)^2+(-1%-0.75%)^2)/SQRT(3)=1.5%